# Worst-Case Symbolic Constraints Analysis and Generalisation with Large Language Models

### v1

@@
Abstract
Large language models (LLMs) have been successfully applied to a variety of coding tasks, including code generation, completion, and repair. However, more complex symbolic reasoning tasks remain largely unexplored by LLMs. This paper investigates the capacity of LLMs to reason about worst-case executions in programs through symbolic constraints analysis, aiming to connect LLMs and symbolic reasoning approaches. Specifically, we define and address the problem of worst-case symbolic constraints analysis as a measure to assess the comprehension of LLMs. We evaluate the performance of existing LLMs on this novel task and further improve their capabilities through symbolic reasoning-guided fine-tuning, grounded in SMT (Satisfiability Modulo Theories) constraint solving and supported by a specially designed dataset of symbolic constraints. Experimental results show that our solver-aligned model, WARP-1.0-3B, consistently surpasses size-matched and even much larger baselines, demonstrating that a 3B LLM can recover the very constraints that pin down an algorithm’s worst-case behaviour through reinforcement learning methods. These findings suggest that LLMs are capable of engaging in deeper symbolic reasoning, supporting a closer integration between neural network-based learning and formal methods for rigorous program analysis.
1 Introduction
@@
Understanding the worst-case algorithmic complexity of software systems has many practical applications, ranging from compiler optimisations, finding and fixing performance bottlenecks, or improving cybersecurity in the presence of denial-of-service attacks. Profilers are typically used to find performance bottlenecks; however, they are inherently incomplete, as they are limited by the number of test inputs used. Symbolic execution techniques, such as WISE (wise-benchmarks) and SPF-WCA (spf-wca), attempt exhaustive analysis but face severe scalability limitations, due to the exponential growth of possible execution paths.
Meanwhile, large language models (LLMs), exemplified by GPT-4 (openai2024gpt4ocard), have demonstrated remarkable capabilities in a variety of programming tasks, including code generation, completion, and repair (poldrack2023ai; fu2023chatgpt). However, their subsymbolic architecture–based on statistical pattern recognition rather than explicit logical reasoning–is not immediately suited for the kinds of rigorous analysis required in worst-case complexity reasoning or formal software verification.
This gap has catalysed the field of neurosymbolic AI, which reconciles the flexibility of neural inference with the precision of symbolic reasoning. Recent work in the automated theorem proving domain has applied various approaches like reinforcement learning, retrieval-augmented generation and integration of external symbolic tools (often called “hammers”) to uplift the capabilities of LLMs in the neuro-symbolic reasoning dimension (xin2024deepseekproverv15harnessingproofassistant; yang2023leandojotheoremprovingretrievalaugmented; jiang2022thorwieldinghammersintegrate).
@@
Building on this momentum, we introduce Worst-case Asymptotic Reasoner for Path constraints (WARP), a unified neuro-symbolic framework for the synthesis and verification of worst-case path constraints, as illustrated in Figure 1. WARP combines symbolic program analysis, reinforcement learning with solver-guided feedback, and rigorous benchmarking to improve the asymptotic reasoning capabilities of LLMs.
Our work advances neurosymbolic reasoning for program analysis by addressing the challenge of integrating formal verification with LLMs, and it makes the following contributions:
Neurosymbolic Framework for Worst-Case Complexity Analysis
@@
At the core of the WARP framework is a novel learning paradigm, which we term Reinforcement Learning from Solver-Guided Feedback (RLSGF). This method fine-tunes LLMs to generate symbolic constraints by leveraging feedback from SMT (Satisfiability Modulo Theories) solvers, for the synthesis of semantically valid and generalisable SMT formulas for analysing software programs. Rather than exhaustively exploring an exponentially growing search space, WARP features training on tractable, small-scale program input constraints, enabling the model to extrapolate constraint patterns to larger inputs—where traditional tools often fail due to path explosion.
@@
To advance and evaluate worst-case–constraint synthesis, we introduce two complementary resources. First, a fine-tuning dataset of synthetically generated programs and their worst-case path constraints is designed to elevate the symbolic-reasoning skills of LLMs. Secondly, a first-of-its-kind benchmark that assesses the semantic fidelity of generated constraints, formal correctness, and generalisation. By focusing on the under-explored task of LLM-driven synthesis of worst-case constraints, the benchmark provides a rigorous, reproducible testbed for future work in neurosymbolic reasoning for worst-case complexity analysis spectrum.
@@ A Trained Model for Worst-Case Symbolic Constraint Reasoning
We develop and release the WARP model, an LLM fine-tuned using the framework to produce symbolic constraints that generalise across input sizes and outperform baselines on asymptotic prediction tasks.
All the data and models for WARP have been made publicly available111
https://figshare.com/articles/online_resource/invaR1ant_RL_Dataset/28743047
https://figshare.com/articles/online_resource/invaR1ant-veRL/28724531
https://figshare.com/articles/online_resource/invaR1ant-benchmark/28724534.
2 Related Work
Symbolic Program Analysis
@@ 2 Related Work
LLMs for Program Analysis
@@
The emergence of LLMs pretrained on code (e.g. GPT-4, Codex) has opened new avenues for program reasoning. Modern LLMs can generate syntactically correct code and even perform simple reasoning about program behaviour in-context. This has spurred research into using LLMs for software verification and analysis tasks beyond code synthesis. For instance, pmlr-v202-pei23a investigated whether LLMs can infer loop invariants and other program properties important for verification. Their study demonstrated that a sufficiently powerful LLM can propose relevant invariants for synthetic Java programs, hinting at the model’s reasoning ability. Similarly, tihanyi2024newerasoftwaresecurity showed that LLMs can assist in vulnerability repair by generating candidate patches which are then validated with formal verification tools. Despite these promising results, purely non-symbolic approaches do not offer formal correctness guarantees. That is, an LLM’s output (e.g., an invariant or a code patch) might seem plausible while being erroneous. Consequently, because sub-symbolic LLMs operate as probabilistic next-token predictors (brown2020languagemodelsfewshotlearners), it is necessary to deploy an external checker (e.g., an SMT solver or verifier) to validate these outputs.
Neurosymbolic Reasoning
@@
Increasingly, researchers are exploring neurosymbolic methods that integrate neural networks with symbolic reasoning to capitalise on the strengths of each. In the context of program verification, neurosymbolic systems use learning to guide or supplement traditional formal analyses. GPT-f (polu2020generativelanguagemodelingautomated) was an early landmark, integrating a transformer-based language model into the Metamath (metamath) formal system as a proof search agent. GPT-f demonstrated that training using the prover’s success to update a value function guiding search can continuously improve theorem-proving performance. Subsequent systems have expanded on this idea. For example, zombori2021findinglongerproofs applied deep reinforcement learning to guide a theorem prover, employing graph neural networks to represent logical formulae, with the aim of handling long, complex proofs. Techniques inspired by AlphaZero’s self-play have also been adapted to propose proof steps and receive verification feedback at each step. While effective, these RL approaches can be computationally expensive due to the need to attempt many proof paths and compute rewards from successful proofs.
@@
An alternative direction has been to use pre-trained LLMs with post-training methods. One notable example is LeanDojo (yang2023leandojotheoremprovingretrievalaugmented), an open-source playground that enables LLMs to interact with the Lean theorem prover (lean_theorem_proving). LeanDojo provides an interface and dataset of proofs, allowing an LLM-based agent to propose proof steps and query the proof assistant. By augmenting an LLM with a retrieval mechanism over a Math library, LeanDojo can select relevant premises and tactics, significantly improving success rates in formal proofs compared to untutored LLM reasoning. DeepSeek Prover (xin2024deepseekproverv15harnessingproofassistant) pushes this synergy further by incorporating reinforcement learning signals from a proof assistant during training. DeepSeek uses Lean to provide a reward signal to fine-tune an LLM for theorem proving. In doing so, the neural model learns to align its generation with formal proof requirements, blending learned intuition with strict verification. Such efforts highlight that combining LLMs with symbolic evaluators (proof assistants, static analysers, solvers) can yield systems that are both creative and more reliable. Our work falls into this neurosymbolic spectrum, where we study and evaluate the application of LLM to synthesise worst-case complexity path constraints in SMT-LIB v2 (SMT2) (BarFT-SMTLIB) format. We further formally verify that the synthesised constraints are semantically equivalent to the worst-case path constraints using the SMT solver Z3 (z3-demoura). The verification is used to guide training and evaluate model capability.
3 Problem Formulation
@@
Symbolic worst-case analysis (WCA) identifies input conditions which software exhibits its maximal resource usage – whether time, memory or instruction count. Identifying these conditions is crucial to prevent vulnerabilities and bottlenecks in real-world applications. However, existing symbolic execution tools, such as SPF-WCA, face significant scalability limitations due to exponential path growth.
Formally, symbolic execution explores the program paths by solving the path constraints generated from conditional branches. The worst-case complexity emerges for input size from the paths yielding the maximal resource consumption where is the cost function for path .
Motivating Example
@@
Consider the example QuickSort (see LABEL:appendix:motivating_example). Worst-case analysis can derive constraints for small inputs (e.g. ) succinctly represented as SMT2 formula:
@@
As input size grows (e.g. ), the symbolic execution tool’s computation becomes increasingly intensive, quickly exceeding thousands of logical clauses. In particular, this represents a -fold increase in the number of execution paths compared to input above. Critically, the complexity of the symbolic execution process itself grows exponentially with input size: .
@@
Even with optimised symbolic methods, like history-preserving guidance, this quickly becomes computationally infeasible at larger input sizes, requiring paradoxical amounts of computational resources and time. Analysing QuickSort at entails exploring paths, translating to impractical CPU time and memory usage, making exhaustive methods unscalable.
Our goal is to generalise from the small-scale worst-case constraints to infer valid worst-case path conditions for larger input sizes, where symbolic execution becomes computationally infeasible. This would allow us to get insights into the worst-case program behavior (and possibly generate exploits) for large input sizes that are reflective of realistic deployment setting.
Augmenting Symbolic Execution with LLMs
To address the inherent scalability limitations of symbolic execution, we introduce the use of LLMs. LLMs, known for their generalisation and semantic reasoning abilities, offer a novel way to extrapolate symbolic constraints. Specifically, we task the LLM to: (1) Learn patterns from constraints derived at tractable input sizes; (2) Extrapolate these patterns to effectively predict symbolic constraints at significantly larger scales, thereby avoiding the need for and limitations of exhaustive symbolic exploration at larger input sizes.
4 Methodology
@@
Following the workflow illustrated in Figure 1, we introduce WARP-1.0-3B, an LLM fine-tuned through reinforcement learning to synthesise SMT2 constraints for neurosymbolic tasks. Alongside the model, we present a benchmark designed to evaluate constraint generalisation. Our results show that WARP-1.0-3B matches the performance of larger proprietary and open-source baselines in terms of accuracy and semantic validity of generated constraints, highlighting its effectiveness as a lightweight neurosymbolic alternative. This work not only advances neurosymbolic reasoning but also serves as the groundwork for future research in automated constraint verification.
4.1 Worst-case Path Constraints Curation
In order to accommodate the novel application of LLMs for asymptotic prediction of worst-case path constraints, a corpus of Java programs and the worst-case execution path constraints that follow have been collected and formatted to enable the development of generalisation and predictive capabilities from LLMs through reinforcement learning.

### v2

@@
Abstract.
Large language models (LLMs) have demonstrated strong performance on coding tasks such as generation, completion and repair, but their ability to handle complex symbolic reasoning over code still remains underexplored. We introduce the task of worst-case symbolic constraints analysis, which requires inferring the symbolic constraints that characterise worst-case program executions; these constraints can be solved to obtain inputs that expose performance bottlenecks or denial-of-service vulnerabilities in software systems. We show that even state-of-the-art LLMs (e.g., GPT-5) struggle when applied directly on this task. To address this challenge, we propose WARP, an innovative neurosymbolic approach that computes worst-case constraints on smaller concrete input sizes using existing program analysis tools, and then leverages LLMs to generalise these constraints to larger input sizes.
Concretely, WARP comprises: (1) an incremental strategy for LLM-based worst-case reasoning, (2) a solver-aligned neurosymbolic framework that integrates reinforcement learning with SMT (Satisfiability Modulo Theories) solving, and (3) a curated dataset of symbolic constraints. Experimental results show that WARP consistently improves performance on worst-case constraint reasoning. Leveraging the curated constraint dataset, we use reinforcement learning to fine-tune a model, WARP-1.0-3B, which significantly outperforms size-matched and even larger baselines. These results demonstrate that incremental constraint reasoning enhances LLMs’ ability to handle symbolic reasoning and highlight the potential for deeper integration between neural learning and formal methods in rigorous program analysis.
1. Introduction
@@
Understanding the worst-case algorithmic complexity of software systems has many practical applications, ranging from compiler optimisations, finding and fixing performance bottlenecks, or improving cybersecurity in the presence of denial-of-service attacks. Profilers are typically used to find performance bottlenecks; however, they are inherently incomplete, as they are limited by the number of test inputs used. Symbolic execution techniques, such as WISE (wise-benchmarks) and SPF-WCA (spf-wca), attempt exhaustive analysis but face severe scalability limitations, due to the exponential growth of possible execution paths. Meanwhile, large language models (LLMs), exemplified by GPT-5 (openai-gpt5-systemcard), have demonstrated remarkable capabilities in a variety of programming tasks, including code generation, completion, and repair (poldrack2023ai; fu2023chatgpt). However, their sub-symbolic architecture–based on statistical pattern recognition rather than explicit logical reasoning–is not immediately suited for the kinds of rigorous analysis required in worst-case complexity reasoning or formal software verification. This gap has catalysed the field of neurosymbolic AI, which reconciles the flexibility of neural inference with the precision of symbolic reasoning. Recent work in the automated theorem proving domain has applied various approaches like reinforcement learning, retrieval-augmented generation and integration of external symbolic tools (often called “hammers”) to uplift the capabilities of LLMs in the neurosymbolic reasoning dimension (xin2024deepseekproverv15harnessingproofassistant; yang2023leandojotheoremprovingretrievalaugmented; jiang2022thorwieldinghammersintegrate).
@@
Building on this momentum, we introduce Worst-case Asymptotic Reasoner for Path constraints (WARP), a unified neurosymbolic framework for the synthesis and verification of worst-case path constraints, as illustrated in Figure 1. WARP combines symbolic program analysis, reinforcement learning with solver-guided feedback, and rigorous benchmarking to improve the asymptotic reasoning capabilities of LLMs. Our work advances neurosymbolic reasoning for program analysis by addressing the challenge of integrating formal verification with LLMs, and it makes the following contributions:
Incremental Worst-Case Constraint Reasoning
As one of our primary contributions, we introduce an incremental inference strategy for worst-case constraint reasoning with LLMs: obtain solver-checked worst-case constraints for small input sizes using existing analysis tools, then (train and) prompt LLMs to generalise these constraint patterns to larger inputs. This reframes the task as semantics-preserving pattern induction, sharply reduces the effective search space, mitigates path explosion, and yields substantial gains over direct prompting on large inputs.
Neurosymbolic Approach for Worst-Case Complexity Analysis
@@
At the core of the WARP framework is a novel learning paradigm, which we term Reinforcement Learning from Solver-Guided Feedback (RLSGF). This method fine-tunes LLMs to generate symbolic constraints by leveraging feedback from SMT solvers, for the synthesis of semantically valid and generalisable SMT formulas for analysing software programs. Rather than exhaustively exploring an exponentially growing search space, WARP features training on tractable, small-scale program input constraints, enabling the model to extrapolate constraint patterns to larger inputs—where traditional tools often fail due to path explosion.
@@
To advance and evaluate worst-case–constraint synthesis, we introduce two complementary resources. First, a fine-tuning dataset of synthetically generated programs and their worst-case path constraints is designed to elevate the symbolic-reasoning skills of LLMs. Secondly, a first-of-its-kind benchmark that assesses the semantic fidelity of generated constraints, formal correctness, and generalisation. By focusing on the under-explored task of LLM-driven synthesis of worst-case constraints, the benchmark provides a rigorous, reproducible testbed for future work in neurosymbolic reasoning for WCA spectrum.
@@ A Trained Model for Worst-Case Symbolic Constraint Reasoning
We develop and also release WARP-1.0-3B, an LLM fine-tuned using the framework to produce symbolic constraints that generalise across input sizes and outperform baselines on asymptotic prediction tasks.
2. Related Work
2.1. Symbolic Program Analysis
@@
2.2. LLMs for Program Analysis
@@
The emergence of LLMs pretrained on code (e.g. GPT-4, Codex) has opened new avenues for program reasoning. Modern LLMs can generate syntactically correct code and even perform simple reasoning about program behaviour in-context. This has spurred research into using LLMs for software verification and analysis tasks beyond code synthesis. For instance, (pmlr-v202-pei23a) investigated whether LLMs can infer loop invariants and other program properties important for verification. Their study demonstrated that a sufficiently powerful LLM can propose relevant invariants for synthetic Java programs, hinting at the model’s reasoning ability. Similarly, (tihanyi2024newerasoftwaresecurity) showed that LLMs can assist in vulnerability repair by generating candidate patches which are then validated with formal verification tools. Despite these promising results, purely non-symbolic approaches do not offer formal correctness guarantees. That is, an LLM’s output (e.g., an invariant or a code patch) might seem plausible while being erroneous. Consequently, because sub-symbolic LLMs operate as probabilistic next-token predictors (brown2020languagemodelsfewshotlearners), it is necessary to deploy an external checker (e.g., an SMT solver or verifier) to validate these outputs.
2.3. Neurosymbolic Reasoning
@@
Increasingly, researchers are exploring neurosymbolic methods that integrate neural networks with symbolic reasoning to capitalise on the strengths of each. In the context of program verification, neurosymbolic systems use learning to guide or supplement traditional formal analyses. GPT-f (polu2020generativelanguagemodelingautomated) was an early landmark, integrating a transformer-based language model into the Metamath (metamath) formal system as a proof search agent. GPT-f demonstrated that training using the prover’s success to update a value function guiding search can continuously improve theorem-proving performance. Subsequent systems have expanded on this idea. For example, (zombori2021findinglongerproofs) applied deep reinforcement learning to guide a theorem prover, employing graph neural networks to represent logical formulae, with the aim of handling long, complex proofs. Techniques inspired by AlphaZero’s self-play have also been adapted to propose proof steps and receive verification feedback at each step. While effective, these RL approaches can be computationally expensive due to the need to attempt many proof paths and compute rewards from successful proofs.
@@
An alternative direction has been to use pre-trained LLMs with post-training methods. One notable example is LeanDojo (yang2023leandojotheoremprovingretrievalaugmented), an open-source playground that enables LLMs to interact with the Lean theorem prover (lean_theorem_proving). LeanDojo provides an interface and dataset of proofs, allowing an LLM-based agent to propose proof steps and query the proof assistant. By augmenting an LLM with a retrieval mechanism over a Math library, LeanDojo can select relevant premises and tactics, significantly improving success rates in formal proofs compared to untutored LLM reasoning. DeepSeek Prover (xin2024deepseekproverv15harnessingproofassistant) pushes this synergy further by incorporating reinforcement learning signals from a proof assistant during training. DeepSeek uses Lean to provide a reward signal to fine-tune an LLM for theorem proving. In doing so, the neural model learns to align its generation with formal proof requirements, blending learned intuition with strict verification. Such efforts highlight that combining LLMs with symbolic evaluators (proof assistants, static analysers, solvers) can yield systems that are both creative and more reliable. Our work falls into this neurosymbolic spectrum, where we study and evaluate the application of LLM to synthesise worst-case complexity path constraints in SMT-LIB v2 (SMT) (BarFT-SMTLIB) format. We further formally verify that the synthesised constraints are semantically equivalent to the worst-case path constraints using the SMT solver Z3 (z3-demoura). The verification is used to guide training and evaluate model capability.
3. Problem Formulation
@@
Symbolic worst-case analysis identifies input conditions under which software exhibits its maximal resource usage – whether time, memory, or instruction count. Identifying these conditions is crucial to prevent vulnerabilities and bottlenecks in real-world applications. However, existing symbolic execution tools, such as SPF-WCA, face significant scalability limitations due to the exponential path growth. Symbolic execution explores the program paths by solving the path constraints generated from conditional branches. The worst-case complexity emerges for input size from the paths yielding the maximal resource consumption where is the cost function for path .
3.1. Motivating Example
@@
Consider the example QuickSort (see 1). Worst-case analysis can derive constraints for small inputs (e.g. ) succinctly represented as SMT formula:
@@
As input size grows (e.g. ), the symbolic execution tool’s computation becomes increasingly intensive, quickly exceeding thousands of logical clauses. In particular, this represents a -fold increase in the number of execution paths compared to input above.
@@
Even with optimised symbolic methods, like history-preserving guidance, this quickly becomes computationally infeasible at larger input sizes, requiring paradoxical amounts of computational resources and time. Analysing QuickSort at entails exploring paths, translating to impractical CPU time and memory usage, rendering exhaustive methods computationally infeasible.
| WARP Benchmark Programs | Input Size | ||||
| 2 | 4 | 8 | 16 | 30 | |
| ArrayTwister | ✗ | ✗ | ✗ | ✗ | ✗ |
| BinarySearch | ✗ | ✓ | ✓ | ✓ | ✗ |
| BinarySearchTreeHeight | ✓ | ✗ | ✓ | ✓ | ✓ |
| BinaryTreeSearch | ✗ | ✗ | ✗ | ✗ | ✗ |
| BubbleSort | ✓ | ✓ | ✓ | ✗ | ✗ |
| CaseFlipper | ✗ | ✗ | ✗ | ✗ | ✗ |
| Collatz | ✗ | ✗ | ✗ | ✗ | ✗ |
| ComplexStateMachineParser | ✗ | ✗ | ✗ | ✗ | ✗ |
| Dijkstra | ✗ | ✗ | ✗ | ✗ | ✗ |
| DizzyRamp | ✗ | ✗ | ✗ | ✗ | ✗ |
| GreedyStepper | ✗ | ✗ | ✗ | ✗ | ✗ |
| KnapsackSolver | ✗ | ✗ | ✗ | ✗ | ✗ |
| MazeSolver | ✗ | ✗ | ✗ | ✗ | ✗ |
| MergeSort | ✗ | ✗ | ✗ | ✗ | ✗ |
| NaiveFibonacci | ✗ | ✗ | ✗ | ✗ | ✗ |
| QuickSort | ✓ | ✗ | ✗ | ✗ | ✗ |
| RampUp | ✗ | ✗ | ✗ | ✗ | ✗ |
| SortedListInsert | ✗ | ✗ | ✗ | ✗ | ✗ |
| SubarraySumFinder | ✓ | ✗ | ✗ | ✓ | ✗ |
| TowerOfHanoi | ✗ | ✗ | ✗ | ✗ | ✗ |
Notes: ✓= TRUE, ✗= FALSE
3.2. Vanilla Prompting
We first explore the direct application of an LLM to generate SMT worst-case constraints for the Java programs. The LLM is given the program and it is instructed to output a syntactically correct, canonical SMT assertion (see 2). To quantify performance under vanilla prompting, we evaluated a state-of-the-art proprietary LLM, GPT-5 (openai-gpt5-systemcard), across all programs in the WARP benchmark (see subsection 4.1). For each program, we sampled five input sizes and tasked the model with generating the worst-case path constraints at each size. A generated constraint is counted as correct only if it is both: (1) syntactically valid and (2) semantically equivalent to the ground-truth worst-case constraint verified by the Z3 solver. This dual criterion ensures that correctness reflects true logical equivalence rather than superficial string similarity.
As summarised in Table 1, GPT-5 achieves only in overall accuracy. Most outputs are syntactically well-formed but fail semantic equivalence checks, underscoring the difficulty of worst-case constraint reasoning: without additional structure, even current state-of-the-art models cannot reliably generate correct constraints. These findings motivate the incremental neurosymbolic reasoning strategy of WARP: rather than requiring the model to induce full constraints from scratch, we scaffold learning with solver-verified instances at smaller input sizes, enabling systematic extrapolation to larger and otherwise infeasible regimes.
3.3. Augmenting Symbolic Execution with LLMs
We propose a neurosymbolic approach that uses SPF-WCA and LLMs in an incremental fashion. Concretely, we use SPF-WCA to obtain constraints at small input sizes and we task the LLM to:
Learn constraint patterns from those tractable input sizes.
Extrapolate these patterns to predict constraints for significantly larger inputs.
Definition 3.1 (Incremental Worst-Case Constraint Reasoning).
Given a sequence of solver-derived constraint sets for small input sizes , infer a constraint set for a larger input size such that with respect to solver-based semantic equivalence.
This incremental formulation reframes worst-case analysis as a pattern induction problem: instead of exploring all paths for large , the learner generalises from small-scale constraints to approximate or reproduce . This reduces the exponential search space, mitigates path explosion, and enables analysis at input scales where symbolic execution alone is infeasible.
4. Methodology
@@
Figure 1 gives a high-level overview of WARP. Note that WARP includes a benchmark designed to evaluate constraint generalisation, and it is applied to finetune an LLM through reinforcement learning to synthesise SMT constraints for our neurosymbolic task. Our results show that the incremental reasoning approach significantly improves LLMs’ ability to solve worst-case constraint reasoning tasks. The fine-tuned WARP-1.0-3B model also matches the performance of much larger proprietary and open-source baselines in terms of both accuracy and semantic validity of generated constraints. Therefore, our results underscore WARP’s effectiveness as a lightweight neurosymbolic alternative.
4.1. Incremental Worst-Case Path Constraints Curation
The defining feature of WARP is its incremental approach to worst-case constraint reasoning. Instead of asking an LLM to derive large- constraints from scratch, we first curate solver-verified constraints at small input sizes where symbolic execution is efficient and reliable. These small- constraints act as building blocks: they capture the recurring structural patterns of worst-case program behaviour and provide the context needed for the LLM to extrapolate to larger, harder cases. Concretely, we use SPF-WCA (spf-wca) to analyse programs at tractable input sizes in its default configuration. SPF-WCA’s default cost model utilises the depth of the symbolic execution path, i.e., the more branches with symbolic conditions are visited during the execution, the more costly such a path is. For each program and size , SPF-WCA yields worst-case path constraints in SMT format. By collecting these constraints across multiple small- values, we build sequences that illustrate how the constraint structure evolves as increases.
WARP leverages these curated constraints in an incremental fashion during both inference and training. During inference, curated examples at smaller values are included in the prompt, and the model is tasked with generating the constraint for a larger, held-out input size. During fine-tuning, they form the foundation of a dataset that familiarises LLMs with the syntax and semantics of SMT constraints while teaching them the progression of constraint patterns across input sizes. This setup explicitly isolates the ability to generalise incrementally, ensuring that observed performance gains reflect learned symbolic reasoning patterns rather than access to ground-truth large- solutions.
4.2. Dataset Creation and Benchmarking